OpenStudy (ny,ny):

The altitude of an object in feet is given by h=-16t^2+8t+48. What is its maximum altitude? After how many seconds 't' will it strike the ground?

OpenStudy (anonymous):

ok so what your going to use is the quadratic formula \[-b \pm \sqrt{(b)^{2}-4ac}\div2a\] so first plug in the equation

OpenStudy (anonymous):

after plunging in the equation tell me what you got

OpenStudy (ny,ny):

x=-13.61 or x=14.11 approximately.

OpenStudy (anonymous):

i am not sure about that can you quickly show your work?

OpenStudy (ny,ny):

\[-8\pm \sqrt{-8^{2}-4(-16)(48)}\]

OpenStudy (ny,ny):

and then divided by -32

OpenStudy (anonymous):

h=-16t^2+8t+48 h=-16t^2+8t-1+1+48 h= 49 - (4t-1)^2 thus max altitude is 49 when t=1/4

OpenStudy (anonymous):

by putting h=0 one can find after how many sec it reache s the ground thus 49 - (4t-1)^2=0 (4t-1)^2=49 4t-1=+-7 t =(1+-7)/4 since t cannot be negative hence reqd t=(1+7)/4=2

OpenStudy (anonymous):

@ NY,NY check the alternate method shown

OpenStudy (ny,ny):

thanks. :)

OpenStudy (anonymous):

welcome dear